OpenStudy (anonymous):

shelly is 3 years older than Michele. four years ago the sum of their age was 67. find the age of each person now

OpenStudy (anonymous):

hmm if 4 years ago the sum of their ages was 67, what is the sum of their ages now?

OpenStudy (istim):

Shelly = 3 + M 67 = (Shelly + M) Sub in Shelly. 67 = (3 + M + M) 67 = 3 + 2M 64 = 2M 32 = M Remember, this was 4 years ago, so add 4 Michelle = 36 Now recall, that Shelly is actually 3 years OLDER than Michelle. Shelly = 39.